### v2

# Precise Mobile Manipulation of Small Everyday Objects

Abstract
Many everyday mobile manipulation tasks require precise interaction with small objects, such as grasping a knob to open a cabinet or pressing a light switch. In this paper, we develop Servoing with Vision Models (SVM), a closed-loop framework that enables a mobile manipulator to tackle such precise tasks involving the manipulation of small objects. SVM uses state-of-the-art vision foundation models to generate 3D targets for visual servoing to enable diverse tasks in novel environments. Naively doing so fails because of occlusion by the end-effector. SVM mitigates this using vision models that out-paint the end-effector thereby significantly enhancing target localization. We demonstrate that aided by out-painting methods, open-vocabulary object detectors can serve as a drop-in module for SVM to seek semantic targets (e.g. knobs) and point tracking methods can help SVM reliably pursue interaction sites indicated by user clicks. We conduct a large-scale evaluation spanning experiments in 10 novel environments across 6 buildings including 72 different object instances. SVM obtains a 71% zero-shot success rate on manipulating unseen objects in novel environments in the real world, outperforming an open-loop control method by an absolute 42% and an imitation learning baseline trained on 1000+ demonstrations also by an absolute success rate of 50%.
Figure 1: Many everyday mobile manipulation tasks require reaching a precise interaction site before executing a motion primitive, e.g. precise reaching of a knob / handle to pull open a cupboard in (a) and (b), or precisely reaching a user-indicated button / book before pushing it in (c) and (d) (shown via the red dot). Open loop execution is unable to meet the high-precision needed for these tasks. In this paper, we develop Servoing with Vision Models (SVM), a training-free framework that closes the loop to enable a commodity mobile manipulator to tackle these tasks.
I Introduction
Figure 2: When using off-the-shelf detectors on wrist camera data, knob detections (indicated by the red point in the second column) are incorrect. Errors stem from occlusion of the knob due to the end-effector (top) and due to the presence of the end-effector (out-of-distribution object) even when the knob is unoccluded (bottom). Out-painting the end-effector (right two columns) fixes this.
Mobile manipulators hold the promise of performing a wide range of useful tasks in our everyday environments. However, a major obstacle to realizing this vision lies in the lack of precise mobile manipulation capabilities of current systems. Many real world tasks require precise interaction with small objects, such as grasping a knob to pull open a cabinet or pressing a light switch, where even a small deviation can cause failure. A mobile manipulator’s mobility makes these tasks even more challenging. For example, small errors during navigation, say on a thick carpeted surface, can easily exceed the tight tolerance required for precise tasks, and a non-holonomic base may limit precise repositioning. Furthermore, mobility means that mobile manipulators have to manipulate in varied locations under diverse lighting and unnatural viewpoints (e.g. looking top-down at a drawer from very close by), demanding significantly broader generalization than stationary manipulators confined to a fixed environment. As a result, developing mobile manipulators capable of performing precise tasks while generalizing to diverse environments remains an open problem.
Many precise mobile manipulation tasks involve stylized interactions with small objects: reaching a precise interaction site before executing a simple motion. The difficulty in these tasks lies in reaching the accurate pre-task pose around the small target object whereas the subsequent motion is easy to execute. For getting to the accurate pre-task pose, open-loop execution using a sense-plan-act paradigm does not work because inaccuracies in perception and control prevent correct engagement with the small interaction site. This necessitates a closed-loop approach. Imitation learning would be a natural choice, but as our experiments will show, policies trained with imitation fail to achieve sufficient precision and generalization even when trained on 1000+ real world demonstrations. So, how can we achieve precise mobile manipulation of small everyday objects?
Visual servoing with an eye-in-hand camera is an effective technique to close the loop to precisely pursue targets [1] without requiring large-scale task-specific training for broad generalization, unlike closed-loop imitation learning. However, vanilla visual servoing makes strong assumptions, e.g. requiring known 3D objects or target images, which are not available in our in-the-wild setting. Our proposed approach, Servoing with Vision Models or SVM, marries together visual servoing with modern vision foundation models to mitigate these limitations. This leads to an effective system which is able to operate in a closed-loop manner, and at the same time is versatile enough to operate in novel environments on previously unseen objects.
SVM leverages vision foundation models in two ways. First, we use them to specify targets for the visual servoing module. This alleviates the need for known 3D objects or target images. We experiment with two ways to specify targets: a) semantic categories, and b) points of interaction. For objects that have a well-known semantic category (e.g. drawer knobs or cabinet handles), we use an open-world object detector (e.g. Detic [2]) to continuously detect the target during visual servoing. However, not all mobile manipulation interaction sites, e.g. the different buttons on a microwave, correspond to a semantic category. We tackle such cases by using point trackers (e.g. CoTracker [3]) to continuously track a user-indicated interaction site (e.g. a single user click in the first image, specifying which button on the microwave to push) during visual servoing. Thus, the use of vision foundation models takes care of the target specification problem in visual servoing.
Figure 3: Servoing with Vision Models (SVM) is a framework for precise reaching for mobile manipulators. Starting from an input RGB-D wrist camera image with a target specified either via a semantic label (e.g. handle) or a user-clicked point on the image, SVM outputs whole-body control commands to convey the end-effector to the target location by closing the loop with visual feedback. SVM first paints out the end-effector using a video outpainting model, uses vision foundation models to continuously detect the target object (or track the desired target point) to compute 3D servoing targets, which are passed to a servo to obtain whole-body control commands (see Section IV).
One problem, however, still remains. The use of visual servoing with an eye-in-hand camera for manipulation tasks suffers due to occlusion of the environment by the manipulator. Such occlusion can be particularly detrimental if it leads to out-of-distribution input to the vision model that now starts producing erroneous predictions (see second column of Figure 2). We mitigate this issue using yet another advance in computer vision: video inpainting models [4, 5, 6]. We out-paint the robot end-effector to obtain a clean view of the scene (see right two columns of Figure 2). This improves the detection performance of the vision foundation model, leading to improved overall success.
We conduct experiments in 10 novel environments across 6 buildings to test SVM on several real world tasks: grasping a knob to pull open drawers / cupboards, grasping a handle to pull open a cabinet, pushing buttons, toggling switches, and pushing books into place on bookshelves. We obtain a 71% success rate on these challenging tasks, zero-shot on novel objects in previously unseen environments. As expected, SVM performs much better than open-loop control which only succeeds on 29% of trials. SVM without end-effector out-painting only obtains a 62% success rate, suggesting that the presence of the end-effector does degrade performance and that our out-painting strategy is able to successfully mitigate it. Somewhat surprisingly, vision foundation models (Detic [2], CoTracker [3], and ProPainter [5] in our experiments) perform quite well on out-of-distribution wrist camera images.
Perhaps most interesting is the comparison of SVM vs. imitation learning. As discussed, fine-grained, closed-loop tasks are precisely where one would expect imitation learning to excel over a modular learning approach. So it is natural to ask how well SVM performs vs. imitation learning. We conduct this comparison on the tasks of pulling knobs and handles for opening articulated objects. Specifically, we compare to the recent Robot Utility Models (RUM) work from Etukuru et al. [7]. RUM is a closed-loop policy trained on 1200 demonstrations for opening cabinets and 525 demonstrations for opening drawers and thus serves as a very strong imitation learning comparison point. Surprisingly, we find that SVM outperforms RUM by an absolute success rate of 50% across $30$ trials.
To summarize, we develop SVM, an approach to enable precise mobile manipulation in the real world. This is made possible by marrying together visual servoing with modern vision foundation models. Our experiments reveal the effectiveness of our proposed approach over a number of strong baselines. These results suggest that SVM can serve as an effective alternative to imitation learning for generalizable and precise mobile manipulation of small everyday objects.
II Related Work
Vision Foundation Models (Detection, Point Tracking, and Inpainting). Training on Internet-scale datasets [8, 9, 10] with large-capacity models [11] has dramatically improved the generalization performance of vision systems. This coupled with alignment of visual representations with ones from language (e.g. CLIP [8]) has led to effective open-vocabulary object detectors, e.g. Detic [2], OVR-CNN [12]. Similar advances in diffusion-based generative models [13, 14, 15, 16] and large-scale training have led to effective image generation models. These models have been leveraged for image and video inpainting [5, 4]. Inpainting models have also been used in robotics to mitigate domain gap between human and robot data [17, 4]. Last, point-based tracking in videos is seeing renewed interest in recent times [18, 3, 19]. Given a set of 2D points in the first frame, these models are able to track them over a video. Use of machine learning makes these new approaches more robust than earlier versions [20]. Forecasts of point tracks into the future have been used as a intermediate representation for policy learning in robotics [21, 22].
Visual Servoing. Visual servoing (image-based, pose-based, and hybrid approaches) outputs control commands that convey the camera (and the attached manipulator) to a desired location [23, 1, 24]. Research has investigated use of different features to compute distance between current and target images: photometric distance [25], matching histograms [26], features from pre-trained neural networks [27], and has even trained neural networks to directly predict the relative geometric transformation between images [28]. Visual servoing has been applied for manipulation [29], navigation [30, 31, 32], 1-shot visual imitation [33] and for seeking far away targets via intermediate view synthesis [34].
Precise Manipulation. Many recent works have tackled various precise manipulation tasks [35, 36, 37]. ForceSight [38] uses a neural network to predict visual-force goals for precise manipulation tasks on the Stretch. However, it trains a specialized affordance model on a small dataset of objects, in contrast to our approach of leveraging a vision foundation model, which subsequently leads to ForceSight’s poor generalization (see supplementary video for more details). Most similar to our work is MOSART [39], which tackles opening of articulated objects using a modular system. Despite its strong generalization performance, this is an open-loop method, and as our experiments will show, struggles on high-precision tasks. We introduce a novel methodology for closing-the-loop for such modular systems to enable precise manipulation.
Imitation Learning (IL). IL [40, 41] is a general tool for learning closed-loop manipulation policies. Recent research has made many advances: expressive policy architectures [42, 43], output representations [44], and use of object-centric representations [45]. However, this generality comes with the need for a large number of demonstrations for generalization [46]. IL has also been successfully applied to eye-in-hand settings [47, 48, 7]. Recent one/few-shot imitation learning methods [49, 50, 51] leverage the structure of the task (getting to a bottleneck pose + motion replay) to learn from a one / few demonstrations but are then restricted to interacting with the object they were trained on. We also leverage the same structure in tasks, but by employing vision foundation models trained on large datasets, we are able to operate on novel objects in novel environments.
Figure 4: Our evaluation consists of 10 environments across 6 buildings, including 72 different object instances. Note that we exclusively test on novel objects in novel environments not used for training or development in any manner.
III Task
Many everyday household tasks involve precise manipulation followed by execution of a motion primitive. Examples include grasping a knob or a handle to pull open drawers / cupboards, or pushing a button on a microwave. We consider two variants—interaction sites specified via a semantic label (e.g. knob or handle) or via a user-defined point (e.g. a click on an image specifying the button to push)—for a total of six tasks, and assume that the motion primitive is provided or easily specified. Our goal is to enable a low DoF commodity mobile manipulator equipped with a RGB-D wrist camera to accomplish such high precision tasks in previously unseen environments.
IV Method
At a high level, SVM uses visual servoing on eye-in-hand images to guide the end-effector to the interaction site. One of the key innovations is leveraging vision foundation models to detect or track the target for visual feedback. Additionally, to handle occlusion from the end-effector, we first paint it out before utilizing the vision foundation models. Let $I_{t}$ be the wrist image and let $x_{t}$ be the robot configuration. The output actions are computed as follows:
$a_{t}=\pi\left(x_{t},g\left(f\left(I_{t},[I_{1},\ldots,I_{t-1}]\right)\right)\right),$ (1)
where $f(I,{\bf I})$ is a video inpainting function that paints out the end-effector from image $I$ using images in $\bf I$} as reference, $g(I)$ localizes the target in 3D in the wrist camera frame, and $\pi$ computes the desired joint velocities using the current robot state $x_{t}$ and the current target location output by $g$. Figure 3 shows an overview of our proposed method and we describe each component below.
Inpainting. Given RGB images from the wrist camera, the inpainting function $f$ uses past frames from the wrist camera to inpaint the current frame $I_{t}$. We utilize a video inpainting method (as opposed to an image inpainting method) for better performance: a video inpainting model has access to previous frames (where the object may be unoccluded), which can lead to improved inpainting. We adopt ProPainter [5], a transformer-based video inpainting model, to realize $f$.
If the target object is occluded in the first frame, even video inpainting may fail to reconstruct it. To address this, we execute a one-time “look-around” primitive that moves the end-effector vertically and laterally to gather scene context. We then use the resulting ten frames as input to the inpainting model, which is limited to processing the ten most recent images to reduce inference time.
Out-painting the end-effector also requires a mask of the end-effector. We use a manually constructed mask that coarsely covers the end-effector. We find this to work better for out-painting than a fine mask of the end-effector obtained using the image segmentation model SAM [10].
Table I: Execution Success Rates. We compare Servoing with Vision Models (SVM) to a previous system (MOSART [39]), open-loop execution using targets computed in the wrist camera, and a version of SVM without inpainting in 10 novel environments across 6 buildings. Tasks require precise control and open-loop execution fails. MOSART’s contact correction works for handles but struggles with knobs and it cannot handle user-clicked targets. Inpainting matters for semantic targets.
Semantic Targets User-clicked Targets Total
Knobs (Cabinets) Knobs (Drawers) Handles (Cabinets) Light Button Book Push Toggle Switch
MOSART [39][i.e. Open Loop (Head Camera)] 1/7 0/7 6/7 - - - 7/42
Open-Loop (Eye-in-Hand) 2/7 0/7 6/7 0/7 4/7 0/7 12/42
SVM w/o inpainting 6/7 3/7 4/7 5/7 4/7 4/7 26/42
SVM (Ours) 6/7 5/7 6/7 5/7 4/7 4/7 30/42
Interaction Site Localization. Given an image with the end-effector painted out, our next goal is to localize the object of interest to obtain the 3D location of the target. We handle the two specifications for the interaction site, via a semantic label or a user click, separately as described below.
  • Detection. For semantically specified targets (e.g. knobs / handles), we use Detic [2], an open-vocabulary detector trained on large-scale datasets. We prompt Detic with the object class ‘handle’ for handles and ‘knob’ for knobs. If Detic detects multiple handles in the image, we select the handle closest to the center of the image. Detic also outputs a mask for the object. We compute the center of the mask and use this as 2D position of the object of interest.
  • Tracking. For tasks specified via a user click, e.g. the point on the book or the button in Figure 1, we make use of CoTracker [3], a point tracking method for videos. Given a point in the first frame, CoTracker is able to track it over subsequent frames seen during execution. CoTracker notes that tracking performance is better when tracking many points together. We therefore sample 40 points randomly around the user click and find this to drastically improve tracking performance.
Both methods yield a 2D target location, which we lift to 3D using the depth image. If the target lies within the end-effector mask (i.e. is occluded), we use the depth from the nearest prior frame.
Closed-loop Control. Given the interaction sites’ 3D location, we employ visual servoing to realize $\pi$ to compute velocity control commands. Visual servoing computes whole-body velocity control commands that minimize the distance between the end-effector and the 3D target point [1].
V Experiments
We conduct experiments in 10 previously unseen environments across 6 previously unseen buildings including 72 different object instances (see Figure 4). Our experiments are designed to test a) the extent to which open-loop execution is an issue for precise mobile manipulation tasks, b) the effectiveness of blind proprioceptive correction techniques, c) whether object detectors and point trackers perform reliably in wrist camera images, d) the impact of end-effector occlusion and the efficacy of video inpainting in mitigating it, and e) how our proposed SVM framework compares to large-scale imitation learning.
V-A Tasks and Experimental Setup
We use the Stretch RE2, a commodity mobile manipulator with a 5DOF arm on a non-holonomic base, upgraded with the Dex Wrist 3 and an eye-in-hand RGB-D camera (Intel D405). We evaluate SVM across 3 task families: a) grasping knobs to open drawers/cabinets, b) grasping handles to open cabinets, and c) pushing objects (buttons, books, switches). Our focus is on generalization. Therefore, we exclusively test on previously unseen instances, not used during development.
All tasks involve precise manipulation, followed by execution of a motion primitive. For the pushing tasks, the precise motion is to get the end-effector exactly at the indicated point and the motion primitive is to push in the direction perpendicular to the surface, before retracting the end-effector upon contact. The robot is positioned such that the target position is within the field of view of the wrist camera. A user selects the point of pushing via a mouse click on the wrist camera image. The goal is to push at the indicated location. Success is determined by whether the push results in the desired outcome (light turns on / off or book gets pushed in). Since the rubber gripper bends on contact, we attach a rigid tool and account for its geometry during servoing.
For the opening articulated object tasks, the precise manipulation involves grasping a knob or handle, followed by whole-body motion to open the cabinet. We adopt MOSART [39] for this motion. Each episode begins with the robot $\approx 1.5m$ from the object, visible from the head camera. MOSART computes articulation parameters and navigates to a pre-grasp pose. SVM (or baseline) then centers the gripper around the knob/handle, after which MOSART resumes: it extends until contact, closes the end-effector, and finally executes the predicted motion plan. Following [39], success is defined as opening the cabinet by over $60^{\circ}$ or the drawer by more than 24cm.
For the precise manipulation part, all baselines consume the current and previous RGB-D images from the wrist camera and output full body motor commands.
V-B Baselines
We compare against four other methods for the precise manipulation part of these tasks.
Figure 5: SVM vs. Open-Loop (Eye-in-Hand) baseline. (top) In opening a cabinet with a knob, slight errors in getting to the target cause the end-effector to slip off, leading to failure for the baseline, whereas our method is able to successfully complete the task. (bottom) Slight errors in getting to the target cause failure, whereas SVM successfully turns the lights off. Note the high quality of CoTracker’s track ( blue dot).
MOSART [39]. The recent modular system for opening cabinets and drawers [39] reports impressive performance with open-loop control (using the head camera from 1.5m away), combined with proprioception-based feedback to compensate for errors in perception and control when interacting with handles. We test if such correction is also sufficient for interacting with knobs. Note that such correction is not possible for the smaller buttons and pliable books.
Open-Loop (Eye-in-Hand). To assess the precision requirements of the tasks and to set it in context with the manipulation capabilities of the robot platform, this baseline uses open-loop execution starting from estimates for the 3D target position from the first wrist camera image.
SVM (no inpainting). To assess the negative impact of the end-effector’s presence in the wrist camera image during manipulation, we ablate the use of inpainting.
Robot Utility Models (RUM) [7]. For the opening articulated object tasks, we also compare to Robot Utility Models (RUM), a closed-loop imitation learning method recently proposed by Etukuru et al. [7]. RUM is trained on a substantial dataset comprising expert demonstrations, including 1,200 instances of cabinet opening and 525 of drawer opening, gathered from roughly 40 different environments. This dataset stands as the most extensive imitation learning dataset for articulated object manipulation to date, establishing RUM as a strong baseline for our evaluation.
For RUM, similar to SVM, we use MOSART to compute articulation parameters and convey the robot to a pre-grasp location with the target handle in view of the wrist camera. One of RUM’s assumptions is a good view of the handle. To benefit RUM, we try out three different heights of the wrist camera, and report the best result for RUM.
V-C Results
Table I and Table II present results from our experiments. We discuss our key experimental findings from our large-scale evaluation spanning real world experiments across 10 novel environments below.
Vision foundation models with servoing enable generalizable and precise mobile manipulation. Our approach SVM successfully solves 71% of unseen object instances in unseen environments that were not used for development.
Vision foundation models work reasonably well even on wrist camera images. Inpainting works well on wrist camera images (see Figure 2). Closing the loop using feedback from vision detectors and point trackers on wrist camera images also work well, particularly when we use inpainted images. See some example detections and point tracks in Figure 5. Detic [2] was able to reliably detect the knobs and handles, and CoTracker [3] was able to successfully track the point of interaction letting us solve 30/42 task instances.
Closing the loop is necessary for these precise tasks. While the proprioception-based strategies proposed in MOSART [39] work out for handles, they are inadequate for targets like knobs and just do not work for tasks like pushing buttons. Using estimates from the wrist camera is better, but open-loop execution still fails for knobs and pushing buttons (see Figure 5).
Erroneous detections without inpainting hamper performance on handles and our end-effector out-painting strategy effectively mitigates it. As shown in Figure 2, presence of the end-effector caused the object detector to misfire, leading to failed execution. Our out-painting approach mitigates this issue, leading to a higher success rate than the approach without out-painting. Interestingly, CoTracker [3] is quite robust to occlusion and does not benefit from inpainting.
Table II: Comparison of SVM vs. RUM [7], a recent large-scale end-to-end imitation learning method trained on 1200 demos for opening cabinets and 525 demos for opening drawers across 40 different environments. Our evaluation spans objects from 9 previously unseen environments across 6 buildings. SVM significantly outperforms RUM. Note that RUM was not trained on knobs.
Handle Knobs Total
Cabinets Drawer Cabinets Drawer
RUM [7] 2/8 4/7 0/7 1/8 7/30
SVM (Ours) 6/8 7/7 4/7 5/8 22/30
Figure 6: SVM vs RUM [7]. We show roll-outs of both SVM and RUM: at the start of the episode, at the start of the closed-loop grasping, just before closing the gripper, and during manipulation.
Figure 7: SVM Failure Modes. We characterize the failure modes of SVM of all the 20 failures we encountered in Table I and Table II. On the right, we show some examples of these failure modes.
Closed-loop imitation learning struggles on novel objects. As presented in Table II, SVM significantly outperforms RUM in a paired evaluation on unseen objects across 9 novel environments spanning 6 buildings. For handles, RUM succeeds on 6/15 handles vs. SVM succeeds on 13/15 handles. Although RUM was not explicitly trained on knobs, its exposure to a wide variety of handles during training suggested potential for generalization, which motivated our evaluation on knobs as well. While RUM achieves a non-zero success rate on knobs, it struggles relative to SVM (see Figure 6). One failure mode of RUM is failing to grasp the handle despite approaching it closely. These results demonstrate that a modular approach that leverages the broad generalization capabilities of vision foundation models is able to generalize much better than end-to-end imitation learning trained on 1000+ demonstrations, which must learn all aspects of the task from scratch.
SVM Failure Modes. Across 20 failures of SVM in both Table I and Table II, 35% were perception errors and 65% were execution errors (see Figure 7). Within perception, mis-detections were surprisingly the dominant failure mode, despite using a vision foundation model. Within execution, insufficient precision was the dominant cause of failure. This suggests that future work should focus on strengthening eye-in-hand perception models and improving precise control under limited degrees of freedom.
VI Discussion
In this paper, we describe SVM, a framework for precise manipulation tasks that involve reaching a precise interaction site followed by execution of a primitive motion. Vision foundation models help us mitigate issues caused by occlusion by the end-effector, thereby enabling the use of off-the-shelf open-vocabulary object detectors and point trackers to estimate servoing targets during execution. The use of strong off-the-shelf models also provides broad generalization for perception, while the use of servoing provides robust control. This enables SVM to solve tasks in novel environments on novel objects, obtaining a 71% zero-shot success rate across 6 precise mobile manipulation tasks. Perhaps most surprisingly, despite being modular, SVM outperforms the strong end-to-end imitation learning system RUM [7] which was trained on over 1000 demonstrations. This is particularly striking as imitation learning is often the tool of choice for precise tasks requiring closed-loop execution.
VII Limitations
Even though SVM performs quite well across many tasks on novel objects in novel environments, it suffers from some shortcomings. Running these large vision models is computationally expensive and we have to offload computation to a A40 GPU sitting in a server. Even with this GPU, we are only able to run the vision pipeline at a 0.1 Hz leading to slow executions. Building vision models specialized to wrist camera images may work better and faster. A second limitation is the reliance on depth from the wrist camera which may be poor in some situations e.g. shiny or dark objects. Use of learned disparity estimators [52] with stereo images could mitigate this. As our focus is on the precise reaching of interaction sites, we work with a hand-crafted task decomposition into the interaction site and primitive motion. In the future, we could obtain such a decomposition using LLMs, or from demonstrations, thereby expanding the set of tasks we can tackle.
VIII Acknowledgements
This material is based upon work supported by DARPA (Machine Common Sense program), an NSF CAREER Award (IIS-2143873), and the Andrew T. Yang Research and Entrepreneurship Award. We are grateful to the Centre for Autonomy for lending us the Stretch RE2 robot used in this work. We thank Aditya Prakash and Xiaoyu Zhang for their feedback on manuscript.